∼ ∼ public

key asymmetric secret key symmentric session key ¨ o SSH Communications Security Datafellows $ client only

The authenticity of host 'nms.tyict.vtc.edu.hk (172.19.64.56)' can't
  be established.
RSA key fingerprint is
  48:2a:cb:1f:ac:7d:21:89:39:de:47:20:cf:06:d3:44
Are you sure you want to continue connecting (yes/no)?
Warning: Permanently added 'nms.tyict.vtc.edu.hk'(RSA) to the list of known hosts.

scp neighbour’s IP:file ∼

nickl
server

scp nick@server:/usr/share/emacs/site-lisp/site-start.el ∼ scp file neighbour’s IP:/tmp/directory option purpose must be kept secret host session user client server much user agent

yes

∼ Diffie-Hillman Key Exchange host keys ssh-keygen -t rsa

$ ssh-keygen -t rsa
Generating public/private rsa key pair.
Enter file in which to save the key (/home/nicku/.ssh/id_rsa):
Enter passphrase (empty for no passphrase):
Enter same passphrase again:
Your identification has been saved in /home/nicku/.ssh/id_rsa.
Your public key has been saved in /home/nicku/.ssh/id_rsa.pub.
The key fingerprint is:
13:6f:6b:1f:ac:6a:21:89:f0:de:47:20:d4:06:5c:2b

∼ ∼ ∼ must only ∼ as one line ∼ ∼ ∼ scp -p ∼/.ssh/id_{r}sa.pubremoteIP : ssh remote IP ∼ mkdir ∼/.sshremote chmod 700 ∼/.ssh ∼ ∼ cat ∼/id_{r}sa.pub >>∼/.ssh/authorized_{k}eysremote rm ∼/id_{r}sa.pubremote exit client server client client not ∼ not ∼ not ∼ client also SSH1 file SSH2 file ∼ ∼ ?? ?? ∼ ∼ ∼ ∼ ∼ ∼

FileMin PermissionMax Permission

∼ ∼ ∼ ∼ ∼ ∼ ∼ refuse ∼ not ∼ is client client ∼ $ Note client $ x.x

$ sudo install -m 755 keychain /usr/bin
keychain ~/.ssh/id_rsa